Use de Moivre's Theorem to derive a formila for the 4th root of -8
1
Expert's answer
2020-10-19T16:46:32-0400

Let "z= \cos \theta +i\sin\theta." Then "z^4= \cos 4\theta +i\sin4\theta." Hence "z^4=-1\Rightarrow \cos 4\theta =-1, \sin 4\theta=0." Hence "4\theta= 2k\pi+\pi." Hence "\theta= k\pi/2 +\pi/4."

Hence "z=\cos (k\pi/2 +\pi/4)+isin(k\pi/2+\pi/4)." Now "sin , cos" being periodic of period "2\pi," "z" has distinct values for "k=0,1,2,3." Hence for 4th root of -8 the solutions are "8^{1/4}[cos(k\pi/2+\pi/4)+i sin(k\pi/2+\pi/4)]"